# Dynamic compensation for pump-induced frequency shift in Kerr-cat qubit initialization

### v2

@@ Abstract
The noise-biased Kerr-cat qubit is an attractive candidate for fault-tolerant quantum computation; however, its initialization faces challenges due to the squeezing pump-induced frequency shift (PIFS). Here, we propose and demonstrate a dynamic compensation method to mitigate the effect of PIFS during the Kerr-cat qubit initialization. Utilizing a novel nonlinearity-engineered triple-loop SQUID device, we realize a stabilized Kerr-cat qubit and validate the advantages of the dynamic compensation method by improving the initialization fidelity from 57% to 78%, with a projected fidelity of 91% after excluding state preparation and measurement errors. Our results not only advance the practical implementation of Kerr-cat qubits, but also provide valuable insights into the fundamental adiabatic dynamics of these systems. This work paves the way for scalable quantum processors that leverage the bias-preserving properties of Kerr-cat qubits.
@@
Two main approaches exist for realizing cat qubits: dissipative cat qubits Mirrahimi et al. (2014); Lescanne et al. (2020) and Kerr-cat qubits Puri et al. (2017); Grimm et al. (2020). Kerr-cat qubits are particularly appealing due to their potential for high-speed gate operations Xu et al. (2022). However, their initialization, which typically involves adiabatically ramping up a two-photon pump to map the Fock states to the cat state manifold Puri et al. (2017); Grimm et al. (2020), faces a significant challenge: the pump-induced frequency shift (PIFS). This shift, arising from either charge Grimm et al. (2020) or flux Lu et al. (2023); Iyama et al. (2024) pumping, can cause decoherence and state leakage during the initialization Grimm et al. (2020); Puri et al. (2020). Previous attempts to compensate for PIFS rely on applying a static detuning between the pump and qubit frequencies Grimm et al. (2020). While this approach works for small PIFS, it fails to account for the dynamic nature of PIFS during the ramping-up process, leading to imperfect initialization. Moreover, as the pursuit of better bias-preserving properties necessitates larger cat amplitudes, the effect of PIFS becomes increasingly pronounced and demands a more sophisticated compensation strategy.
In this work, we introduce a dynamic compensation method that continuously adjusts the pump frequency to match the instantaneous PIFS during the adiabatic initialization. We experimentally demonstrate the advantage of this dynamic compensation for a Kerr-cat qubit using a novel Nonlinearity-Engineered Triple-loop SQUID (NETS) device Hua et al. . This device not only holds promise for realizing bias-preserving CNOT gates Puri et al. (2020); Chamberland et al. (2022), but also offers independent tuning of nonlinearities compared with other Kerr-cat qubit candidates such as SNAIL Frattini et al. (2017, 2018); Sivak et al. (2019), ATS Lescanne et al. (2020), Josephson ring modulator Bergeal et al. (2010); Roch et al. (2012), and a series of asymmetric DC SQUIDs Iyama et al. (2024). Our theoretical and experimental results showcase the superiority of the dynamic compensation approach, particularly for large PIFS values. Our demonstration paves the way for high-fidelity Kerr-cat qubit operations and advances the field towards fault-tolerant quantum computation.
@@
Figure 2: (a) Circuit schematic of the device. The Nonlinearity-Engineered Triple-loop SQUID (NETS) is shunted with a large capacitor, similar to that of a transmon. (b) Circuit schematic of the NETS. $I_{2}$ is the flux pumping port. The loop areas are meticulously designed such that the flux ratios are 1:-1:3, and the ratio of the small (pink) to the large (blue) Josephson junctions is $r\approx 0.16$. More details and parameters about the device are provided in Ref. SM . (c) Photograph of the NETS. (d) Illustration comparing the dynamic and static compensation methods. The envelope $\varepsilon_{2}(t)$ follows a tanh function. In the static method, the pump frequency remains fixed over time, whereas in the dynamic method, the pump frequency varies with time $\Delta_{\mathrm{as}}(t)\propto\varepsilon^{2}_{2}(t)$. (e-l) Comparison of the ideal and experimental Wigner functions for two typical PIFS. (e-h) The first two even-parity eigenstates for $|\alpha|^{2}\approx 1.1$ and 1.9, with $\Delta_{\mathrm{as,KC}}=3.4\,\mathrm{MHz}<K$ and $\Delta_{\mathrm{as,KC}}=9.7\,\mathrm{MHz}>K$. (i-l) The experimental results based on the dynamic and static methods. The color contrast is mainly limited by the decoherence of the transmon.
@@
Experiments.- The Kerr-cat qubit is implemented using a novel NETS device with its schematic circuit shown in Fig. 2(a). An ancillary transmon qubit is introduced for characterizing the Kerr-cat qubit. The NETS consists of three loops [Fig. 2(b)], holding the advantage of independently tuning the nonlinearities of different orders by controlling the external flux through three current lines $I_{1}$, $I_{2}$, and $I_{3}$ Hua et al. ; SM . Figure 2(c) depicts the fabricated NETS device, with the ratios of the junctions and loop areas being meticulously designed. By biasing the NETS at $(\phi_{\mathrm{e1}},\phi_{\mathrm{e2}},\phi_{\mathrm{e3}})=(-0.08,1.08,0)\pi$ and applying a flux pump via $I_{2}$ line, the Hamiltonian of Eq. 1 for the Kerr-cat qubit can be realized SM .
Without an external pump, the NETS functions as a nonlinear oscillator at a frequency of $\omega_{\mathrm{a}}/2\pi=5.6\,\mathrm{GHz}$ with a Kerr nonlinearity $K/2\pi=6.9\,\mathrm{MHz}$. The amplitude damping and transverse coherence times are $T_{\mathrm{1N}}=6.0\,\mathrm{\mu s}$ and $T_{\mathrm{2N}}=3.0\,\mathrm{\mu s}$, respectively. The Kerr-cat qubit is realized by applying a flux pump at $\omega_{\mathrm{s}}/2\pi\approx 11.2\,\mathrm{GHz}$ with an amplitude of $\varepsilon_{\mathrm{2,KC}}/2\pi=9.7\,\mathrm{MHz}$ to the NETS. By initializing the NETS in the state $(\left|0\right\rangle+\left|1\right\rangle)/\sqrt{2}$, the Kerr-cat qubit can be prepared in the $\left|\alpha\right\rangle$ state on the equator of the Bloch sphere. The precession due to PIFS is measured by varying the holding time of NETS in the Kerr-cat states, allowing the calibration of $\Delta_{\mathrm{as,KC}}$ SM . Furthermore, the Rabi drive of the Kerr-cat qubit is realized by a charge drive on the NETS, with a Rabi frequency along the X-axis given by $\Omega_{X}=\mathrm{Re}(4\epsilon_{\mathrm{x}}\alpha)/\sqrt{1-\exp(-4|\alpha|^{2})}$  Puri et al. (2020). Here, $\epsilon_{\mathrm{x}}$ is the Rabi drive strength. By fitting the Rabi dynamics, the Kerr-cat size is calibrated as $|\alpha|^{2}=1.41$, with a significantly extended damping time of $T_{\mathrm{1,KC}}=13\,\mathrm{\mu s}$ SM .
@@
The fidelities of the mapping based on the two schemes are further characterized by implementing a symmetric ramping-down process to remap the Kerr-cat states back to the Fock space. This whole process is ideally an identity channel, and can be directly characterized by performing quantum process tomography (QPT) Nielsen and Chuang (2010) in the Fock state basis when the pump is off. Figure 3(a) presents the process fidelity for different pump amplitudes, i.e., different PIFS and $|\alpha|^{2}$, obtained using the two schemes. It demonstrates that the dynamic compensation method consistently outperforms the static compensation method across the entire range of PIFS values. In particular, when $\Delta_{\mathrm{as,KC}}$ is near $K$ or $3K$, the static compensation method shows strong oscillations of the fidelity. This can be explained by the collapse of the energy levels of the Fock states with respect to the compensated detunings, and the condition for the adiabatic evolution is violated. This leads to the observation of Landau-Zener transitions Landau (1932); Zener and Fowler (1932): an initial Fock state can be mapped to two stabilized states of the pumped NETS when the energy gap between the Fock state and other higher levels becomes small. Due to the energy difference of the stabilized states, the accumulated phase leads to an interference between the two states when mapping the states back to the Fock state basis, exhibiting the sharp oscillations. The evolutions of the energy levels corresponding to these specific cases are shown in Figs. 3(b) and 3(c), respectively. It is worth noting that the squeezing operation preserves parity, which explains the absence of transitions around $\Delta_{\mathrm{as,KC}}=2K$. These observations are also validated by numerical simulations based on QuTip Johansson et al. (2012, 2013), showing excellent agreement between the experimental and numerical results.
The static compensation method shows comparable performance to the dynamic compensation approach when $K<\Delta_{\mathrm{as,KC}}<3K$. This can be attributed to the fact that a precisely symmetric ramping-down process can adiabatically map the stabilized NETS state back to the original Fock space, regardless of whether the NETS state is inside or outside the Kerr-cat qubit space. To directly characterize the initialization of Kerr-cat qubits, we perform a direct cat readout Grimm et al. (2020); Ding et al. (2024); SM of the NETS by coherently converting its state to the readout cavity. Measurements along different axes are realized with an X/2 gate implemented through Rabi drives and a Z/2 gate achieved by instantaneously switching off the flux pump Grimm et al. (2020); SM .
@@
Figure 4(a) shows the experimental QPT results (solid markers) for different $\Delta_{\mathrm{as,KC}}$, confirming several key insights into the performance of the dynamic and static methods. When $\Delta_{\mathrm{as,KC}}/K<1$, both methods exhibit comparable initialization fidelities due to the negligible leakage outside the Kerr-cat qubit space. However, as the PIFS increases beyond $K$, the static compensation method experiences a rapid decline in fidelity due to the leakage. In contrast, the dynamic compensation method maintains a high initialization fidelity throughout the entire range of PIFS values. The simulations incorporating the state preparation and measurement (SPAM) errors (solid lines, mainly contributed by the Z/2 gate SM ) closely match the experimental data, confirming our understanding of the system dynamics and the source of imperfections. Therefore, based on a theoretical model, we can project the achievable fidelity by excluding SPAM errors (dashed line), demonstrating the robustness and effectiveness of the dynamic compensation approach. In particular, for $\Delta_{\mathrm{as,KC}}/K=1.4$ and $|\alpha|^{2}=1.9$, the experimental fidelities are 57% and 78% for these two approaches, respectively, corresponding to 52% and 91% when excluding the SPAM errors. Here, the leakage error of the static compensation method is underestimated due to the imperfections in detecting the leakage states Gautier et al. (2022); Ding et al. (2024). We find that the fidelity degrades with $\Delta_{\mathrm{as,KC}}$ due to the increasing decoherence rates with the Kerr-cat size, and better performance is possible by improving the coherence of the NETS. For instance, a Kerr-cat qubit initialization with a fidelity exceeding 99% is anticipated when the NETS has $T_{\mathrm{1N}}=60\,\mathrm{\mu}$s and $T_{\mathrm{2N}}=30\,\mathrm{\mu}$s.
Conclusion.- We have demonstrated high-fidelity initialization of a Kerr-cat qubit using a novel NETS device with a dynamic compensation method. By carefully analyzing eigenstate evolution of the system under adiabatic pump ramping, we clarify the imperfections arising from PIFS and Landau-Zener transitions. Our experimental and theoretical results consistently show that the dynamic compensation approach outperforms the static compensation scheme, significantly suppressing imperfections associated with large PIFS. At a PIFS to Kerr nonlinearity ratio of $1.4$, the measured initialization fidelity improves from 57% to 78%, indicating a projected fidelity of 91% after excluding SPAM errors. Our demonstration with NETS and dynamic compensation paves the way for future advancements in scalable Kerr-cat qubits, enabling their application in quantum information processing and exploration of noise-biased fault-tolerant techniques.
@@ Acknowledgements.
This work was funded by the National Natural Science Foundation of China (Grants No. 11925404, 92165209, 92365301, 12061131011, 92265210, 11890704, 92365206), Innovation Program for Quantum Science and Technology (Grant No. 2021ZD0300203 and 2021ZD0301800), and the National Key R&D Program (2017YFA0304303). This work was also supported by the Fundamental Research Funds for the Central Universities and USTC Research Funds of the Double First-Class Initiative. This work was partially carried out at the USTC Center for Micro and Nanoscale Research and Fabrication.

### v3

@@ Abstract
The noise-biased Kerr-cat qubit is an attractive candidate for fault-tolerant quantum computation; however, its initialization faces challenges due to the squeezing pump-induced frequency shift (PIFS). Here, we propose and demonstrate a dynamic compensation method to mitigate the effect of PIFS during the Kerr-cat qubit initialization. Utilizing a novel nonlinearity-engineered multi-loop SQUID device, we realize a stabilized Kerr-cat qubit and validate the advantages of the dynamic compensation method by improving the initialization fidelity from 57% to 78%, with a projected fidelity of 91% after excluding state preparation and measurement errors. Our results not only advance the practical implementation of Kerr-cat qubits, but also provide valuable insights into the fundamental adiabatic dynamics of these systems. This work paves the way for scalable quantum processors that leverage the bias-preserving properties of Kerr-cat qubits.
@@
Two main approaches exist for realizing cat qubits: dissipative cat qubits Mirrahimi et al. (2014); Lescanne et al. (2020) and Kerr-cat qubits Puri et al. (2017); Grimm et al. (2020). Kerr-cat qubits are particularly appealing due to their potential for high-speed gate operations Xu et al. (2022). However, their initialization, which typically involves adiabatically ramping up a two-photon pump to map the Fock states to the cat state manifold Puri et al. (2017); Grimm et al. (2020), faces a significant challenge: the pump-induced frequency shift (PIFS). This shift, arising from either charge Grimm et al. (2020) or flux Lu et al. (2023); Iyama et al. (2024); Hoshi et al. (2024) pumping, can cause decoherence and state leakage during the initialization Grimm et al. (2020); Puri et al. (2020). Previous attempts to compensate for PIFS rely on applying a static detuning between the pump and qubit frequencies Grimm et al. (2020). While this approach works for small PIFS, it fails to account for the dynamic nature of PIFS during the ramping-up process, leading to imperfect initialization. Moreover, as the pursuit of better bias-preserving properties necessitates larger cat amplitudes, the effect of PIFS becomes increasingly pronounced and demands a more sophisticated compensation strategy.
In this work, we introduce a dynamic compensation method that continuously adjusts the pump frequency to match the instantaneous PIFS during the adiabatic initialization. We experimentally demonstrate the advantage of this dynamic compensation for a Kerr-cat qubit using a novel Nonlinearity-Engineered Multi-loop SQUID (NEMS) device Hua et al. (2024). This device not only holds promise for realizing bias-preserving CNOT gates Puri et al. (2020); Chamberland et al. (2022), but also offers independent tuning of nonlinearities compared with other Kerr-cat qubit candidates such as SNAIL Frattini et al. (2017, 2018); Sivak et al. (2019); He et al. (2023), ATS Lescanne et al. (2020), Josephson ring modulator Bergeal et al. (2010); Roch et al. (2012), and a series of asymmetric DC SQUIDs Iyama et al. (2024); Hoshi et al. (2024). Our theoretical and experimental results showcase the superiority of the dynamic compensation approach, particularly for large PIFS values. Our demonstration paves the way for high-fidelity Kerr-cat qubit operations and advances the field towards fault-tolerant quantum computation.
@@
Figure 2: (a) Circuit schematic of the device. The Nonlinearity-Engineered Multi-loop SQUID (NEMS) is shunted with a large capacitor, similar to that of a transmon. (b) Circuit schematic of the NEMS. $I_{2}$ is the flux pumping port. The loop areas are meticulously designed such that the flux ratios are 1:-1:3, and the ratio of the small (pink) to the large (blue) Josephson junctions is $r\approx 0.16$. More details and parameters about the device are provided in Ref. SM . (c) Photograph of the NEMS. (d) Illustration comparing the dynamic and static compensation methods. The envelope $\varepsilon_{2}(t)$ follows a tanh function. In the static method, the pump frequency remains fixed over time, whereas in the dynamic method, the pump frequency varies with time $\Delta_{\mathrm{as}}(t)\propto\varepsilon^{2}_{2}(t)$. (e-l) Comparison of the ideal and experimental Wigner functions for two typical PIFS. (e-h) The first two even-parity eigenstates for $|\alpha|^{2}\approx 1.1$ and 1.9, with $\Delta_{\mathrm{as,KC}}=3.4\,\mathrm{MHz}<K$ and $\Delta_{\mathrm{as,KC}}=9.7\,\mathrm{MHz}>K$. (i-l) The experimental results based on the dynamic and static methods. The color contrast is mainly limited by the decoherence of the transmon.
@@
Experiments.- The Kerr-cat qubit is implemented using a novel NEMS device with its schematic circuit shown in Fig. 2(a). An ancillary transmon qubit is introduced for characterizing the Kerr-cat qubit. The NEMS consists of three loops [Fig. 2(b)], holding the advantage of independently tuning the nonlinearities of different orders by controlling the external flux through three current lines $I_{1}$, $I_{2}$, and $I_{3}$ Hua et al. (2024); SM . Figure 2(c) depicts the fabricated NEMS device, with the ratios of the junctions and loop areas being meticulously designed. By biasing the NEMS at $(\phi_{\mathrm{e1}},\phi_{\mathrm{e2}},\phi_{\mathrm{e3}})=(-0.08,1.08,0)\pi$ and applying a flux pump via $I_{2}$ line, the Hamiltonian of Eq. 1 for the Kerr-cat qubit can be realized SM .
Without an external pump, the NEMS functions as a nonlinear oscillator at a frequency of $\omega_{\mathrm{a}}/2\pi=5.6\,\mathrm{GHz}$ with a Kerr nonlinearity $K/2\pi=6.9\,\mathrm{MHz}$. The amplitude damping and transverse coherence times are $T_{\mathrm{1N}}=6.0\,\mathrm{\mu s}$ and $T_{\mathrm{2N}}=3.0\,\mathrm{\mu s}$, respectively. The Kerr-cat qubit is realized by applying a flux pump at $\omega_{\mathrm{s}}/2\pi\approx 11.2\,\mathrm{GHz}$ with an amplitude of $\varepsilon_{\mathrm{2,KC}}/2\pi=9.7\,\mathrm{MHz}$ to the NEMS. By initializing the NEMS in the state $(\left|0\right\rangle+\left|1\right\rangle)/\sqrt{2}$, the Kerr-cat qubit can be prepared in the $\left|\alpha\right\rangle$ state on the equator of the Bloch sphere. The precession due to PIFS is measured by varying the holding time of NEMS in the Kerr-cat states, allowing the calibration of $\Delta_{\mathrm{as,KC}}$ SM . Furthermore, the Rabi drive of the Kerr-cat qubit is realized by a charge drive on the NEMS, with a Rabi frequency along the X-axis given by $\Omega_{X}=\mathrm{Re}(4\epsilon_{\mathrm{x}}\alpha)/\sqrt{1-\exp(-4|\alpha|^{2})}$  Puri et al. (2020). Here, $\epsilon_{\mathrm{x}}$ is the Rabi drive strength. By fitting the Rabi dynamics, the Kerr-cat size is calibrated as $|\alpha|^{2}=1.41$, with a significantly extended damping time of $T_{\mathrm{1,KC}}=13\,\mathrm{\mu s}$ SM .
@@
The fidelities of the mapping based on the two schemes are further characterized by implementing a symmetric ramping-down process to remap the Kerr-cat states back to the Fock space. This whole process is ideally an identity channel, and can be directly characterized by performing quantum process tomography (QPT) Nielsen and Chuang (2010) in the Fock state basis when the pump is off. Figure 3(a) presents the process fidelity for different pump amplitudes, i.e., different PIFS and $|\alpha|^{2}$, obtained using the two schemes. It demonstrates that the dynamic compensation method consistently outperforms the static compensation method across the entire range of PIFS values. In particular, when $\Delta_{\mathrm{as,KC}}$ is near $K$ or $3K$, the static compensation method shows strong oscillations of the fidelity. This can be explained by the collapse of the energy levels of the Fock states with respect to the compensated detunings, and the condition for the adiabatic evolution is violated. This leads to the observation of Landau-Zener transitions Landau (1932); Zener and Fowler (1932): an initial Fock state can be mapped to two stabilized states of the pumped NEMS when the energy gap between the Fock state and other higher levels becomes small. Due to the energy difference of the stabilized states, the accumulated phase leads to an interference between the two states when mapping the states back to the Fock state basis, exhibiting the sharp oscillations. The evolutions of the energy levels corresponding to these specific cases are shown in Figs. 3(b) and 3(c), respectively. It is worth noting that the squeezing operation preserves parity, which explains the absence of transitions around $\Delta_{\mathrm{as,KC}}=2K$. These observations are also validated by numerical simulations based on QuTip Johansson et al. (2012, 2013), showing excellent agreement between the experimental and numerical results.
The static compensation method shows comparable performance to the dynamic compensation approach when $K<\Delta_{\mathrm{as,KC}}<3K$. This can be attributed to the fact that a precisely symmetric ramping-down process can adiabatically map the stabilized NEMS state back to the original Fock space, regardless of whether the NEMS state is inside or outside the Kerr-cat qubit space. To directly characterize the initialization of Kerr-cat qubits, we perform a direct cat readout Grimm et al. (2020); Ding et al. (2024); SM of the NEMS by coherently converting its state to the readout cavity. Measurements along different axes are realized with an X/2 gate implemented through Rabi drives and a Z/2 gate achieved by instantaneously switching off the flux pump Grimm et al. (2020); SM .
@@
Figure 4(a) shows the experimental QPT results (solid markers) for different $\Delta_{\mathrm{as,KC}}$, confirming several key insights into the performance of the dynamic and static methods. When $\Delta_{\mathrm{as,KC}}/K<1$, both methods exhibit comparable initialization fidelities due to the negligible leakage outside the Kerr-cat qubit space. However, as the PIFS increases beyond $K$, the static compensation method experiences a rapid decline in fidelity due to the leakage. In contrast, the dynamic compensation method maintains a high initialization fidelity throughout the entire range of PIFS values. The simulations incorporating the state preparation and measurement (SPAM) errors (solid lines, mainly contributed by the Z/2 gate SM ) closely match the experimental data, confirming our understanding of the system dynamics and the source of imperfections. Therefore, based on a theoretical model, we can project the achievable fidelity by excluding SPAM errors (dashed line), demonstrating the robustness and effectiveness of the dynamic compensation approach. In particular, for $\Delta_{\mathrm{as,KC}}/K=1.4$ and $|\alpha|^{2}=1.9$, the experimental fidelities are 57% and 78% for these two approaches, respectively, corresponding to 52% and 91% when excluding the SPAM errors. Here, the leakage error of the static compensation method is underestimated due to the imperfections in detecting the leakage states Gautier et al. (2022); Ding et al. (2024). We find that the fidelity degrades with $\Delta_{\mathrm{as,KC}}$ due to the increasing decoherence rates with the Kerr-cat size, and better performance is possible by improving the coherence of the NEMS. For instance, a Kerr-cat qubit initialization with a fidelity exceeding 99% is anticipated when the NEMS has $T_{\mathrm{1N}}=60\,\mathrm{\mu}$s and $T_{\mathrm{2N}}=30\,\mathrm{\mu}$s.
Conclusion.- We have demonstrated high-fidelity initialization of a Kerr-cat qubit using a novel NEMS device with a dynamic compensation method. By carefully analyzing eigenstate evolution of the system under adiabatic pump ramping, we clarify the imperfections arising from PIFS and Landau-Zener transitions. Our experimental and theoretical results consistently show that the dynamic compensation approach outperforms the static compensation scheme, significantly suppressing imperfections associated with large PIFS. At a PIFS to Kerr nonlinearity ratio of $1.4$, the measured initialization fidelity improves from 57% to 78%, indicating a projected fidelity of 91% after excluding SPAM errors. Our demonstration with NEMS and dynamic compensation paves the way for future advancements in scalable Kerr-cat qubits, enabling their application in quantum information processing and exploration of noise-biased fault-tolerant techniques.
@@ Acknowledgements.
This work was funded by the National Natural Science Foundation of China (Grants No. 11925404, 92165209, 92365301, 12061131011, 92265210, 11890704, 92365206), Innovation Program for Quantum Science and Technology (Grant No. 2021ZD0300203, 2021ZD0301800, and 2021ZD0300201), and the National Key R&D Program (2017YFA0304303). This work was also supported by the Fundamental Research Funds for the Central Universities and USTC Research Funds of the Double First-Class Initiative. This work was partially carried out at the USTC Center for Micro and Nanoscale Research and Fabrication.
Note added.- When finalizing this work, we became aware of a related study Iyama et al. (2024) that also utilized a chirped pulse for initialization. While that study focuses on the characterization and dynamics of cat states, our work delves into the physical mechanism of the adiabatic evolution of the state mapping process.